How do I determine the equation of the parabola that has vertex at the origin, axis of symmetry is the x-axis, and passes through the point (-9, 2)?
Answer by lwsshak3(11628)

How do I determine the equation of the parabola that has vertex at the origin, axis of symmetry is the x-axis, and passes through the point (-9, 2)?
***
Standard form of equation:x=A(y-k)^2+h, (h,k)=(x,y) coordinates of the vertex
For given parabola: (h,k)=(0,0)
x=Ay^2
solve for A using coordinates of given point (-9,2)
-9=A*2^2=4A
A=-9/4
equation:x=-9y^2/4